OpenStudy (faithspring):

Write the equation of the line in slope intercept form that goes through the point (7, 2) and has a slope of 4.

OpenStudy (trojanpoem):

y = mx + b m is slope, b is y-intercept y = 4x + b 2 = 14 + b b = -12 y = 4x - 12

OpenStudy (faithspring):

show your work

OpenStudy (faithspring):

how did you get 14 though? @TrojanPoem

OpenStudy (trojanpoem):

I used the point given ( x, y) when x = 7 , y =2